Sarah spent a total of $10 on oranges and apples at the supermarket. If she paid $3 less for oranges than she did for apples, how much did Sarah spend on oranges?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52915)

.

oranges = x    dollars

apples = x+3   dollars


An equation is


    Oranges + Apples = 10  dollars,  or

       x    + (x+3)  = 10

       2x            = 10-3

       2x            = 7

        x            = 7/2 = 3.5


ANSWER.  Sara spent $3.50 for oranges.

Solved.



Answer by greenestamps(13216)



x = dollars spent on oranges
x+3 = dollars spent on apples ($3 more than what was spent on oranges)

The total spent was $10:

(x)+(x+3) = 10

Solve using basic algebra.

Assuming you are working on this problem because you are studying algebra, you should understand how to set up and solve the problem that way, or in a similar formal way.

But you can solve the problem informally, and get some good mental exercise, by solving the problem with logical reasoning and simple mental arithmetic.